# Importance of accounting for student identities and intersectionality for creating equitable and inclusive physics learning environments

### v1

@@ III.3.1 Positive interactions and perceptions
When asked about the gender dynamics in her courses, Esme brings up her academic advisor, who she had positive experiences with: “I’m very comfortable with her, because she’s my advisor here and I’ve had her as a professor many times…she wants us to be comfortable with the people that we interact with. I know we’re about to do a project and she’s trying to make sure not to put the women in our class as the only woman in a group, I know she’s working on that and I definitely am grateful for that, because I definitely would not want to be in a group by myself as a woman with men…” She notes that at least her advisor puts in effort to make sure students feel comfortable with peers that they interact with. Also, this instructor is aware of not having a woman as the only woman in a group [79], which Esme appreciates.
@@
When asked about her overall college experience, Ruby says: “I think I’m very fortunate to go to school at [my college] where the instructors really want you to succeed and I felt so supported in physics, math, and honestly all of my classes here.” She feels like her instructors are encouraging and supporting her academically. This is a good sign that Ruby has positive experiences with her instructors. However, this does not specifically mean that she only has had good experiences with all instructors.
@@
Both Ruby and Francis mention positive interactions with their instructors during office hours. Ruby explains that she had difficulty grasping certain concepts in her courses, such as air resistance, but her professors are very helpful: “I went to my professor about it, and she was so so helpful…” In Ruby’s experience, her professor normalized struggle and helped her realize that not all students must understand a physics concept at the same rate. She also says that she feels comfortable to ask professors questions during office hours and that it is a welcoming environment. Francis explains that she doesn’t attend office hours, unless she has a specific question. However, when she has gone, she has had positive experiences. She says: “it was definitely helpful. So it was after an exam I had a question [about] something I did wrong, so I went and she helped me explain why it was wrong and then guided me through the correct answer, I think it was definitely helpful.” It appears that this particular professor provides good support during office hours and did a good job at answering her questions whenever she went.
@@
Francis also explains that one of her favorite class period was one that focused on marginalized students’ identities in physics and their experiences in physics. In this class, students read articles from marginalized people and then they discussed them during class. The discussions focused on how physics needs to change to become more inclusive: “I think, including something like that, at some point in your undergraduate studies is definitely a good thing, trying to make sure, everyone is aware and actively thinking about how you can do better to make things more inclusive and I guess peers just acknowledging that…people are coming from different backgrounds and not assuming that you know more than someone before talking to them.” Francis praises this instructor for incorporating such a discussion into a class and the value of such discussions for all students.
@@
Chloe says that overall, her experience in physics has been good and that the physics community is “welcoming and positive.” She also adds that her instructors “have really made attempts to make their classroom environments better.” She recalls an example of one of her instructors trying to create an inclusive learning environment by asking a student who was trained in such things who, “comes into our class and takes notes on what the [instructor is] doing and what the classic dynamic is and then also comes around and talks to all the students individually and really asked us about everything, like how the classroom environment is, and how we feel, and then they bring it up with the professor.” It seems like this specific instructor cares about how their students are interacting with each other and how they engage with their students. This instructor had someone evaluate their classroom and then provide feedback. This is very useful as it can help instructors understand how students experience their physics classroom environment.
III.3.2 Negative interactions and perceptions
@@
Here we share negative interactions with and perceptions of some of the instructors that these women had. For some women, their negative perceptions stem from lack of action or interventions during negative interactions with peers. This, in particular, is a reflection of the lack of disciplinary action (if male peers’ behavior is not appropriate) or shows lack of efforts in general to create an equitable and inclusive learning environment by instructors.
@@
As part of the interview, when asked if her professors can do more to prevent male students from negatively stereotyping women in physics, Chloe says: “I’m not sure if they noticed it. I think the professors who tried to notice it do [see biases], but if a professor isn’t really trying to actively improve equity in the classroom then it’s easy to miss” To her, it is not apparent that some of the instructors are aware of biases that male students have towards women. According to Chloe, in order to notice these things, instructors have to be, e.g., alert to male students dominating or showing off or de-valuing women and actively improving their classroom environments. This awareness is essential to instructors then making changes to improve the learning environments.
@@
Ruby says that while working with her lab partner who was condescending towards her, one of her instructors intervened and one of them did not. She says, for one class, “The instructor for my class actually did notice that the student was being very difficult to work with and that he wasn’t collaborating with the group, and she I think spoke to him about it.” Although this lecture instructor addressed the student’s behavior, the lab instructor did not. She explains: “But in the lab I don’t know why they didn’t notice, I think, a part of it was, a lot of our experiments were done outside of the physical classroom because, again with COVID, it was hard to sort of build those relationships with professors, so I think that was a part of the reason, but I also think another reason is when you’re submitting work on time and when things are looking fine, a lot of professors don’t go out of their way to look at the problems and a lot of professors, when necessary, will assume [things are fine] or [don’t] want to have to deal with anything like that.” She also offers an additional reason why the lab instructor did not address the student: “because [my lab partner] knew the answers to a lot of things and…he is smart, he’s just horrible to work with and because of that, it’s possible that, that gave him the right, or in their minds, they gave them the right to act how he was acting.” Thus, Ruby suggests that because her male peer appeared smart, that may have justified his behavior to the professor. Thus, Ruby reveals that instructors have a role in addressing students’ behaviors and not being biased by whether the student appears to be smart because these types of student behaviors are tied to the physics culture.
@@
Paulina shares a few experiences when her professor singled her out and assumed that she was falling behind when she was not: “I’ve had maybe one or two experiences, where the instructor has kind of singled me out in the classroom and asked me questions continuously trying to I guess make sure that I’m keeping up with the class, even though there’s been no signs of me struggling with the material.” It is possible that the professor is aware of Paulina’s marginalized identity as a woman of color and overcompensated by asking her questions in a way that she felt stereotyped. In Paulina’s view, this came across as the professor underestimating her, as she also perceives her peers to do. We note that it could be that this instructor had good intentions. However, the physics culture was not equitable and inclusive so it was not possible for Paulina to view the instructor’s intentions and actions as positive, and she perceived this experience as the instructor singling her out as opposed to checking-in on her.
@@
When asked if she’s been in any situations where she had to stand up for herself, Paulina shares experiences in which other students in her classes stood up to the professor: “…someone else has stepped up and [was] like, Hey, this feels like-I feel like I’m not being accurately listened to, or represented in this atmosphere and what professors will do is that a lot of them actually end up being surprised by students speaking up and by students feeling this way, and so I guess after that, they have a conversation and the next class time, for example, [professors will] try to make it known to students Hey, we should be engaging with each other or trying to discretely say…there should be more interactions with individuals that aren’t necessarily White or that are women in STEM and so I think that’s been the extent [of] things usually, when a student speaks up about it, I’ve seen temporary discussions of it with professors and maybe in the classroom setting, but it feels almost as if, a few weeks later, it kind of goes back to what it used to be” According to Paulina, when put in these types of situations in which students call out professors directly, although a professor may have superficially tried to address the issues brought up by students, it is not very impactful or long-lasting. We suspect that the physics culture is so established that when a professor addresses an issue superficially and passively, it does not have any long term effect on the culture since it was not reinforced in any way and things go back to the way they were very soon.
@@
III.4 Reflections on how various aspects of their identities affect physics experiences
@@
Chloe describes herself as introverted and expressed how this part of her personality has made it difficult to navigate group work. She notices that physics is a very social environment saying: “I think personality is the thing that isn’t always talked about like being very social…versus somebody who’s not…most physicists are nerdy and I tend to be more introverted…outside of the classroom. I think personality does create some divides…” She suggests that for someone who is more quiet like her, it may be difficult to try to engage in conversations with social and extroverted physicists.
@@
Francis was one of the few interviewed women who felt she did belong in physics. This could be attributed to her high school experiences as she went to an all-girls high school in the UK. With that in mind, Francis explains that sometimes when there are very few women in a class or as the only woman in a group, “I didn’t really feel like…I didn’t belong there, it was just weird because that wasn’t really something I’ve experienced before being the only woman in a space. I’m not really the kind of person who speaks up a huge amount in class anyway, but yeah I didn’t really know what was weird about it’s just something different.” She does not explicitly say that she feels alienated if there are very few women or she is the only woman in a group with a bunch of men participating. We hypothesize that her being quiet and being from a different country might make her more accepting of different norms she is being exposed to in her classes even though she finds the dominant participation of her male peers “weird” and feels this is a “different” environment compared to what she was used to in her all-girls high school. It is even possible that her male peers were not as condescending to her as they were to some of the other women due to her being from the UK.
@@
Karina explains that sometimes she feels pressure particularly in her classes in which there are not that many women: “When there’s less women, I feel some of that pressure…to like be smart, if you feel like you’re representing woman…So I think I do feel some of that. I think some of it was also just for me as a person. But yeah I think it would be fair to attribute some of that to feeling, to understanding myself as a woman in that space.” Whether or not she is consciously aware that this pressure is partly from the lack of women’s representation in her classes and stereotypes about who can excel in physics, Karina feels it. She may not be fully aware that she is representing women in physics, but she still wants to appear smart and knowledgeable in her physics classes.
@@
Karina confesses she is afraid of not appearing smart. This fear may be stereotype threat, the fear of confirming a negative stereotype about women in physics. “Maybe logically I wouldn’t think that those people would think women are bad at physics, if I asked them to slow down, so maybe if anything, that would be like a subconscious fear because logically that doesn’t really add up to me, but I think subconsciously maybe, yeah.” Although Karina is confident about her own ability, she seems conflicted about not appearing smart in front of her peers: “I think I actually have a lot of confidence in myself and my ability…I don’t think I am too worried about sounding dumb, yeah, which [may sound like] I was saying kind of the opposite before.” She admits that it may appear that she is contradicting her previous statement about being worried about representing women in a positive way. Karina is an example of how conflicted a female student in physics may be about how their identity may shape their experiences.
@@
Karina explains that her privileged identity, i.e., having private education, has cushioned her from feeling isolated. This is particularly interesting because she explains that the further she progressed in her degree, the less women there were. As a woman, this may be challenging to navigate. However, Karina explains that the further she progressed in college: “I was more comfortable in the department, just like I knew all the professors. I knew the space. I had been you know living in [college town] at [college] for longer by the time I was a junior/senior, so I think I didn’t so much feel like isolated or alienated because of that, maybe I would have felt that more, earlier on…I think again, I’ve been pretty cushioned in physics, I haven’t really had to face a whole lot of awful things that have been key to me from being a person in physics, so I think more than anything, it wasn’t a personal experience, being one of a few women, more than like wow this sucks, I wish this field as a whole for the sake of the field and for the sake of everyone had more representation.” Thus, Karina clarifies that her negative feelings, e.g., about appearing smart and feeling like she is representing women in physics, stemmed from not seeing many women in physics and she didn’t particularly have direct negative experiences with her peers based on her identity as a woman.
@@
When asked about if/how her race/ethnicity played a role in her experiences in physics and astronomy, Esme explains that she was one of two Hispanic students in physics at this college and one of two women during her experience at NASA. She reflects on this issue saying, “I think more gender [than race] in my experience has been affected.” This reflection seems consistent with her previous experiences that she describes. Based on her experiences in general, it seems like gender was more salient in her experiences related to physics. Again we hypothesize that this may partly be the case because of her visual characteristics not “manifesting” her ethnicity in the same manner as Paulina.
@@
We note that Paulina constantly emphasizes how her multiple marginalized identities as a woman and person of color were salient in her interactions with peers and even instructors. She attempted to disentangle the two in the context of her interactions with her female peers, i.e., she felt that they judged her differently based on her racial identity (hypothesized by Paulina). However, throughout her interview, both identities were constantly brought up by her when explaining her interactions with others in the physics learning environments in college.
@@
Paulina also mentions that sometimes she works with two students, one being a man of color. She says that they usually consult each other for help when it comes to problem solving. Although she has not had any conversations with him about what his experiences are like in physics, being a man of color, she believes it might be easier for him to seek out help from other students in the classroom. She says, “he feels more comfortable going up to other students if he doesn’t know how to proceed. I guess it becomes a bit easier for him to engage in the classroom with other individuals…it feels as if his first response is, select, seek out other people of color in the classroom and have…conversations about it [with other students].” Although they do work together, she thinks that his identity as a man might be a privilege over her multiple marginalized identities, as she has had difficulty getting other peers to take what she says seriously.
III.5 Supports and Role Models

### v2

@@ IV.3.1 Positive interactions and perceptions
When asked about the gender dynamics in her courses, Esme brings up her academic advisor, who she had positive experiences with. Esme says that she has become very comfortable with her advisor because of her role both as her advisor and instructor. Esme also explains that during class, her advisor “wants us to be comfortable with the people that we interact with. I know we’re about to do a project and she’s trying to make sure not to put the women in our class as the only woman in a group, I know she’s working on that and I definitely am grateful for that, because I definitely would not want to be in a group by myself as a woman with men…” She notes that at least her advisor puts in effort to make sure students feel comfortable with peers that they interact with. It also appears that this instructor is aware of the negative impacts of a woman being the only woman in a group [80].
@@
When asked about her overall college experience, Ruby feels “very fortunate to go to school at [my college] where the instructors really want you to succeed and I felt so supported in physics, math, and honestly all of my classes here.” She feels like her instructors are encouraging and supporting her academically. This is a good sign that Ruby has positive experiences with her instructors. However, this does not specifically mean that she only has had good experiences with all instructors.
@@
Both Ruby and Francis mention positive interactions with their instructors during office hours. Ruby explains that she had difficulty grasping certain concepts in her courses, but whenever she talks to her professors during office hours, they are very helpful. In Ruby’s experience, her professor normalized struggle and helped her realize that not all students must understand a physics concept at the same rate. She also says that she feels comfortable to ask professors questions during office hours and that it is a welcoming environment. Francis explains that she does not attend office hours, unless she has a specific question. However, when she has gone, she has had positive experiences. She says: “it was definitely helpful.” It appears that this particular professor provides good support during office hours and did a good job at answering her questions whenever she went.
@@
Francis also explains that one of her favorite class period was one that focused on marginalized students’ identities in physics and their experiences in physics. In this class, students read articles from marginalized people and then they discussed them during class. The discussions focused on how physics needs to change to become more inclusive. Francis says, “including something like that at some point in your undergraduate studies is definitely a good thing, trying to make sure everyone is aware and actively thinking about how you can do better to make things more inclusive and I guess, peers just acknowledging that…people are coming from different backgrounds and not assuming that you know more than someone before talking to them.” Francis praises this instructor for incorporating such a discussion into a class and the value of such discussions for all students.
@@
Chloe says that overall, her experience in physics has been good and that the physics community is “welcoming and positive.” She also adds that her instructors “have really made attempts to make their classroom environments better.” She recalls an example of one of her instructors trying to create an inclusive learning environment by asking a trained student to take notes on the instructors’ behavior and classroom dynamics. This trained student also talked with the students in the class and asked questions about the classroom environment and how the students feel. Then, they shared this information with the instructor. It seems like this specific instructor cares about how their students are interacting with each other and how they engage with their students. This practice is very useful as it can help instructors understand how students experience their physics classroom environment.
IV.3.2 Negative interactions and perceptions
@@
Here we share negative interactions with and perceptions of some of the instructors that the women had. For some women, their negative perceptions stem from lack of action or interventions during negative interactions with peers. This, in particular, is a reflection of the lack of disciplinary action (if male peers’ behavior is not appropriate) or shows lack of efforts in general to create an equitable and inclusive learning environment by instructors.
@@
As part of the interview, when asked if her professors can do more to prevent male students from negatively stereotyping women in physics, Chloe says: “I’m not sure if they noticed it. I think the professors who tried to notice it do [see biases], but if a professor isn’t really trying to actively improve equity in the classroom then it’s easy to miss.” To her, it is not apparent that some of the instructors are aware of biases that male students have towards women. According to Chloe, in order to notice these things, instructors have to be, e.g., alert to male students dominating or showing off or de-valuing women and actively improving their classroom environments. This awareness is essential to instructors then making changes to improve the learning environments.
@@
Ruby says that while working with her lab partner who was condescending towards her, one of her instructors intervened and one of them did not. She says, for one class, “The instructor for my class actually did notice that the student was being very difficult to work with and that he wasn’t collaborating with the group, and she I think spoke to him about it.” Although this lecture instructor addressed the student’s behavior, the lab instructor did not. She explains: “But in the lab I don’t know why they didn’t notice, I think part of it was, a lot of our experiments were done outside of the physical classroom because, again with COVID, it was hard to sort of build those relationships with professors…” Ruby also explains that when students are submitting assignments on time, these issues might be overlooked: “a lot of professors don’t go out of their way to look at the problems and a lot of professors, when necessary, will assume [things are fine] or [don’t] want to have to deal with anything like that.” Another reason why the lab instructor did not address the student was “because [my lab partner] knew the answers to a lot of things and…he is smart, he’s just horrible to work with and because of that, it’s possible that, that gave him the right… to act how he was acting.” Thus, Ruby suggests that because her male peer appeared smart, that may have justified his behavior to the professor. Thus, Ruby reveals that instructors have a role in addressing students’ behaviors and not being biased by whether the student appears to be smart because these types of student behaviors are tied to the physics culture.
@@
Paulina shares a few experiences when her professor singled her out and assumed that she was falling behind when she was not. She says, “the instructor has kind of singled me out in the classroom and asked me questions continuously trying to I guess make sure that I’m keeping up with the class, even though there’s been no signs of me struggling with the material.” It is possible that the professor is aware of Paulina’s marginalized identity as a woman of color and overcompensated by asking her questions in a way that she felt stereotyped. In Paulina’s view, this came across as the professor underestimating her, as she also perceives her peers to do. We note that it could be that this instructor had good intentions. However, the physics culture was not equitable and inclusive so it was not possible for Paulina to view the instructor’s intentions and actions as positive, and she perceived this experience as the instructor singling her out as opposed to checking-in on her.
@@
When asked if she’s been in any situations where she had to stand up for herself, Paulina shares experiences in which other students in her classes stood up to the professor: “someone else has stepped up and [was] like,… I feel like I’m not being accurately listened to, or represented in this atmosphere.” She also explains that as a result, physics instructors will be surprised and during the next class would address students, saying, “Hey, we should be engaging with each other or trying to discretely say…there should be more interactions with individuals that aren’t necessarily White or that are women in STEM…”’ However, Paulina emphasizes that despite these types of occurrences, the overall feeling is that: “a few weeks later, it kind of goes back to what it used to be.” According to Paulina, when put in these types of situations in which students call out professors directly, although a professor may have superficially tried to address the issues brought up by students, it is not very impactful or long-lasting. We suspect that the physics culture is so established that when a professor addresses an issue superficially and passively, it does not have any long term effect on the culture since it was not reinforced in any way and things go back to the way they were very soon.
@@
IV.4 Reflections on how various aspects of their identities affect physics experiences
@@
Chloe describes herself as introverted and expressed how this part of her personality has made it difficult to navigate group work. She notices that physics is a very social environment saying: “I think personality is the thing that isn’t always talked about like being very social…versus somebody who’s not…most physicists are nerdy and I tend to be more introverted…outside of the classroom. I think personality does create some divides…” Thus, here Chloe shares her idea of a physicist being nerdy and her not relating to these types of identities. She suggests that for someone who is more quiet like her, it may be difficult to try to engage in conversations with social and extroverted physicists.
We note that several women reflected on how their identity as women impacted their experiences in physics. Some women felt that the lack of women in physics spaces negatively impacted them.
@@
Francis was one of the few women who felt she did belong in physics. This could be attributed to her high school experiences as she went to an all-girls high school in the UK. With that in mind, Francis explains that sometimes when there are very few women in a class or as the only woman in a group. She says that despite feeling like she belongs, “…it was just weird because that wasn’t really something I’ve experienced before being the only woman in a space. I’m not really the kind of person who speaks up a huge amount in class anyway, but yeah I didn’t really know what was weird about it’s just something different.” She does not explicitly say that she feels alienated if there are very few women or she is the only woman in a group with a bunch of men participating. We hypothesize that her being quiet and being from a different country might make her more accepting of different norms she is being exposed to in her classes even though she finds the dominant participation of her male peers “weird” and feels this is a “different” environment compared to what she was used to in her all-girls high school. It is even possible that her male peers were not as condescending to her as they were to some of the other women due to her being from the UK.
@@
Karina explains that sometimes she feels pressure to “be smart,” particularly in her classes in which there are not that many women. She says it feels like “…you’re representing women.” Whether or not she is consciously aware that this pressure is partly from the lack of women’s representation in her classes and stereotypes about who can excel in physics, Karina feels it. Here Karina may be alluding to stereotype threat, the fear of confirming a negative stereotype about oneself, i.e., women are not capable of doing or excelling in physics. She may not be fully aware that she is representing women in physics, but she still wants to appear smart and knowledgeable in her physics classes. This may also hinder her learning in that this awareness and resulting pressure takes up cognitive resources that could have been used to learn physics.
@@
Karina confesses she is afraid of not appearing smart. Again, the fear and anxiety may be correlated with stereotype threat. We also see Karina reflect upon these types of thought processes: “Maybe logically I wouldn’t think that those people would think women are bad at physics, if I asked them to slow down, so maybe if anything, that would be like a subconscious fear because logically that doesn’t really add up to me, but I think subconsciously maybe, yeah.” Although Karina is confident about her own ability, she seems conflicted about not appearing smart in front of her peers: “I think I actually have a lot of confidence in myself and my ability…I don’t think I am too worried about sounding dumb, yeah, which [may sound like] I was saying kind of the opposite before.” She admits that it may appear that she is contradicting her previous statement about being worried about representing women in a positive way. Karina is an example of how conflicted a woman physics student may be about how their identity may shape their experiences.
@@
Karina explains that her privileged identity, i.e., having private education, has cushioned her from feeling isolated. This is particularly interesting because she explains that the further she progressed in her degree, the less women there were. As a woman, this may be challenging to navigate. However, Karina explains that the further she progressed in college: “I was more comfortable in the department, just like I knew all the professors. I knew the space. I had been you know living in [college town] at [college] for longer by the time I was a junior/senior, so I think I didn’t so much feel like isolated or alienated because of that, maybe I would have felt that more, earlier on…I think again, I’ve been pretty cushioned in physics, I haven’t really had to face a whole lot of awful things…” Thus, Karina clarifies that her negative feelings, e.g., the pressure to appear smart and represent women in physics, stemmed from not seeing many women in physics and she didn’t particularly have direct negative experiences with her peers based on her identity as a woman. We also see that her familiarity with professors and the physics space helped her navigate these physics experience.
@@
When asked about if/how her race/ethnicity played a role in her experiences in physics and astronomy, Esme explains that she was one of two Hispanic students in physics at this college and one of two women during her experience at NASA. She reflects on this issue saying, “I think more gender [than race] in my experience has been affected.” This reflection seems consistent with her previous experiences that she describes, in which gender was more salient in her experiences related to physics. Again we hypothesize that this may partly be the case because of her visual characteristics not “manifesting” her ethnicity in the same manner as Paulina.
@@
We note that Paulina constantly emphasizes how her multiple marginalized identities as a woman and person of color were salient in her interactions with peers and even instructors. She attempted to disentangle the two in the context of her interactions with her women peers, i.e., she felt that they judged her differently based on her racial identity (hypothesized by Paulina). However, throughout her interview, both identities were constantly brought up by her when explaining her interactions with others in the physics learning environments in college.
@@
Paulina also mentions that sometimes she works with two students, one being a man of color. She says that they usually consult each other for help when it comes to problem solving. Although she has not had any conversations with him about what his experiences are like in physics, being a man of color, she believes it might be easier for him to seek out help from other students in the classroom. She says that he might be more comfortable approaching other students if he needs help. Paulina also speculates that “it becomes a bit easier for him to engage in the classroom with other individuals…it feels as if his first response is, [to] select, seek out other people of color in the classroom and have…conversations about it [with other students].” Although they do work together, she thinks that his identity as a man might be a privilege over her multiple marginalized identities, as she has had difficulty getting other peers to take what she says seriously.
IV.5 Supports and Role Models
In this section, we share who or what has supported and sustained the women and kept them on their path during their undergraduate physics journey. This support may include role models, i.e., people who they looked up to.
IV.5.1 People providing academic support